Find the mean and the median.
–2, -1 2/3, 2/3, 4, 9
Mean = –2, -1 2/3, 2/3, 4, 9 / 5 = 2
Median = 2/3, 4, 9, -1 2/3, -2 = 9
Is this right. Did I do it right

Answer by longjonsilver(2297)

the mean calcualtion is correct

I have no idea whatsoever what you did for the median.

The median needs the numbers written in order, which you already had. You then re-ordered them somehow? Bizarre!

anyway, –2, -1 2/3, 2/3, 4, 9

and the median in the middlemost value... here it lies equidistant from -1 2/3 and 2/3. That number is? (-1 2/3 + 2/3)/2 = -1/2

--> median = -1/2

jon.